# Diffusion Models in Simulation-Based Inference: A Tutorial Review

### v1

@@ 6 Empirical Evaluation
A variety of design choices have been proposed for diffusion models. To assess their practical impact in SBI, we compare specific configurations and derive practical recommendations. For clarity, we adopt a fixed training protocol (Section A.1) and vary only the model family, the parameterization, or the noise schedule, while training models using the noise-prediction loss in (10). All models were trained using BayesFlow 2.0.7 [bayesflow_2023_software].
@@ 6.1 Case Study 1: Low-Dimensional Benchmarks
Our first study presents the most comprehensive evaluation to date of diffusion models as posterior estimators on popular toy benchmark problems [lueckmann2021benchmarking]. This study aims to highlight differences between the numerous design choices discussed in the preceding sections for the easiest types of problems (low-dimensional parameters, low-dimensional observations, easily obtainable high simulation budgets, see Table 2).
@@
  • Diffusion models under different noise schedules, using EDM noise schedules (variance-preserving (VP) or variance-exploding (VE)) with EDM weighting, and diffusion models using a cosine VP schedule with sigmoid weighting;
@@
Figure 10: Comparison of inference performance across design choices on the benchmark by [lueckmann2021benchmarking] (Case Study 1). We report mean and standard deviation of the C2ST, comparing samples from the diffusion model and the ground truth posterior samples for the 10 datasets from the benchmark (C2ST of 0.5 means approximate and reference posterior cannot be distinguished). As a baseline, we report the mean C2ST for the NPE approach in [lueckmann2021benchmarking] trained on a budget of 100,000 simulations, showing how the difficulty across the problems differs.
Our analysis commences with the widely used set of ten benchmark models from [lueckmann2021benchmarking]. These models span a range of low parameter and observation dimensionalities; for detailed descriptions, we refer to the original study by [lueckmann2021benchmarking]. Because simulations are inexpensive, and all problems are low-dimensional, we train each model for 1,000 epochs, generating 30,000 new simulations per epoch. All neural networks are simple MLPs with residual layers without a separate summary network (Section A.1.1). To assess performance, we evaluate each trained model on the ten held-out simulated observations and compare the inferred posteriors with the benchmark “ground-truth” posteriors. We report the classifier two-sample test (C2ST) for a direct comparison with benchmark baselines (see Section A.1 for a definition).
@@
Diffusion models with a variance-preserving EDM noise schedule achieved the best overall performance, closely followed by cosine-based schedules (Figure 10). In contrast, the variance-exploding EDM schedule performed worse on several tasks, often producing higher C2ST scores. Only on the Lotka-Volterra task did it produce the lowest score, whereas other models struggled more with the required posterior contraction.
@@
Across tasks, differences among the $\mathbf{F}$-, $\boldsymbol{v}$-, and $\boldsymbol{\epsilon}$-parameterizations were generally small (Figure 10). EMA smoothing did not yield systematic improvements. In most cases, EMA and non-EMA variants performed similarly.
@@
Flow matching models were competitive on benchmarks with 5 or more parameters but showed a slight overall performance drop compared to diffusion models (Figure 10). Flow matching and the conditional optimal-transport (COT) variant performed comparably, whereas ignoring the condition for computing the OT plan seems to harm accuracy (Figure A.2). Moreover, the power-law noise schedule with $\rho=-0.6$ improved performance in most tasks, similar to previous studies [wildberger2023flow, orsini2025flow].
@@
Across diffusion and flow matching models, the adaptive two-step SDE solver was both the fastest and the most accurate (Figure 10). Performance differences among ODE and SDE solvers were small (Figure A.1). However, annealed Langevin sampling did not reach the accuracy of ODE or SDE solvers, even with 4 times as many steps, and the predictor–corrector scheme also did not improve C2ST scores (Figure A.1). For ODEs, adaptive higher-order methods consistently outperformed Euler: they were faster and achieved lower error (but with only minor differences) by taking larger, fewer steps, with TSIT5 showing a slight advantage over RK45. Step rejection rates in the adaptive methods aligned with trajectory smoothness: flow matching trajectories did not trigger rejections; EDM-based diffusion models triggered fewer than ten; and cosine-schedule models typically caused ten to twenty.
@@
For low-dimensional problems, we recommend to use diffusion models with a variance-preserving EDM noise schedule with $\mathbf{F}$ parameterization. Across problems and samplers, this emerges as the most reliable choice. For some problems that require higher contraction in the posterior, a variance-exploding noise schedule might be beneficial. Moreover, we recommend the usage of adaptive solvers and, in particular, the two-step adaptive SDE solver in light of faster and more accurate inference. Only the consistency models are faster at the cost of potentially reduced accuracy. We discourage the use of annealed Langevin sampling or predictor-corrector sampling, as they are harder to tune and can lead to poor results.
@@ 6.2 Case Study 2: Large-Scale ODE Benchmark
Figure 11: Case Study 2: Comparison of inference performance across design choices. A Visualization of the reactions described by the [beer2014creating] model. The parameters can vary per experimental conditions and some also for repeated experiments, leading to 72 parameters to be estimated. Blue indicates observed states. B Model simulation (with noise). C Best performing model from each family. We also show MCMC as a baseline. All metrics were scaled such that the optimal value would be 1 (only contraction has no optimal value) (Section A.1). D Diffusion models with cosine noise schedules and different parameterizations: $\mathbf{F}$, $\boldsymbol{v}$, and $\boldsymbol{\epsilon}$ using the ODE sampler. E Diffusion models with variance-preserving and exploding EDM noise schedules using the ODE sampler.
@@
To evaluate the capabilities of diffusion models for SBI in actual practice, we consider a mechanistic model from the PEtab collection of ODE models [hass2019benchmark]. As a representative high-dimensional benchmark, we consider the [beer2014creating] model, which represents a quantitative description of the enzymatic reactions involved in indigoidine biosynthesis (Figure 11A). The model includes experimental condition specific parameters for bacterial growth ($\beta$, $\text{Bac}_{\text{max}}$ and time lag $\tau$), conversion of glutamine (Gln) to cyclized glutamine (cGln) ($k_{\text{syn}}$), dimerization of cGln to indigoidine (Ind) ($k_{\text{dim}}$) and degradation of indigoidine ($k_{\text{deg}}$). The model comprises $19$ experimental conditions with more than $\text{dim}(\mathbf{y})=27{,}000$ measurements (Figure 11B) and $\text{dim}(\boldsymbol{\theta})=72$ parameters. It is one of the most data-rich ODE models in the collection, making it particularly suitable for testing the capacity of inference algorithms to handle large parameter spaces (see Section A.1.2 for more details).
@@
Model performance was assessed on 1,000 datasets using scaled calibration error (scaled such that its maximum is 1), contraction, normalized root mean squared error (NRMSE), and C2ST scores, where the classifier is trained on posterior samples together with the log-likelihood as a data summary. For a single scalar ranking across methods, we report the sum of empirical calibration error (ECE) and NRMSE (see Section A.1 for a definition of the metrics).
@@
Flow Matching OT ODE 0.118 (0.002) 0.046 (0.021) 0.706 (0.026) 0.534 1
Uniform ODE 0.119 (0.002) 0.048 (0.023) 0.706 (0.026) 0.536 2
$\rho=-0.6$ ODE 0.131 (0.009) 0.066 (0.028) 0.653 (0.059) 0.529 11
Diffusion Cosine, VP, $\mathbf{v}$ ODE 0.123 (0.003) 0.049 (0.020) 0.701 (0.023) 0.534 3
EDM, VP, $\mathbf{F}$ SDE 0.127 (0.003) 0.047 (0.019) 0.673 (0.029) 0.514 4
Cosine, VP, $\mathbf{F}$ SDE 0.121 (0.002) 0.058 (0.025) 0.703 (0.020) 0.555 5
Cosine, VP, $\mathbf{v}$ SDE 0.121 (0.002) 0.058 (0.024) 0.702 (0.020) 0.542 6
EDM, VP, $\mathbf{F}$ ODE 0.135 (0.003) 0.048 (0.019) 0.647 (0.030) 0.544 7
Cosine, VP, $\mathbf{F}$ ODE 0.127 (0.003) 0.062 (0.025) 0.688 (0.024) 0.532 8
EDM, VE, $\mathbf{F}$ SDE 0.110 (0.002) 0.091 (0.034) 0.844 (0.014) 0.605 12
EDM, VE, $\mathbf{F}$ ODE 0.109 (0.002) 0.144 (0.053) 0.868 (0.012) 0.689 14
Cosine, VP, $\boldsymbol{\epsilon}$ ODE 0.934 (0.012) 0.067 (0.027) 0.263 (0.164) 0.509 17
Cosine, VP, $\boldsymbol{\epsilon}$ SDE 1.051 (0.010) 0.174 (0.102) 0.338 (0.076) 0.537 18
Diffusion (EMA) EDM, VP, $\mathbf{F}$ SDE 0.146 (0.007) 0.049 (0.020) 0.640 (0.057) 0.550 9
EDM, VP, $\mathbf{F}$ ODE 0.145 (0.007) 0.051 (0.020) 0.624 (0.051) 0.520 10
Consistency Discrete ODE 0.119 (0.002) 0.087 (0.041) 0.727 (0.027) 0.540 13
Continuous ODE 0.185 (0.003) 0.159 (0.069) 0.229 (0.038) 0.579 15
MCMC - MCMC 0.124 (0.000) 0.427 (0.217) 1.000 (0.000) 0.889 16
@@ Results
Across both summary networks, several design choices consistently led to improved accuracy and robustness (Figure 11). Flow matching with a uniform noise schedule and its optimal transport variant achieved the strongest overall performance, exhibiting low error and stable behavior. Diffusion models with cosine schedules in the $\mathbf{v}$- and $\mathbf{F}$-parameterizations—using either ODE or SDE samplers—performed similarly well. The variance-preserving EDM schedule also yielded competitive accuracy. In contrast, the variance-exploding EDM schedule produced higher contraction and lower NRMSE but suffered from increased calibration error in this setting (Table 4), aligning with observations from the low-dimensional benchmarks. Moreover, C2ST showed lower sensitivity in this high-dimensional experiment, even when using log-likelihood summaries, suggesting that it does not reliably discriminate between posterior approximations in this regime (Table 4). In particular, when the update from the prior to the posterior was small, as for $\boldsymbol{\epsilon}$-prediction models, C2ST alone is unreliable as an accuracy measure. The relative ranking of inference methods was consistent across both summary networks, indicating that the performance differences are driven by the inference algorithms themselves rather than the choice of summary backbone.
Training times were similar across methods, only flow matching OT took twice as long due to additional computation of the optimal couplings; however, inference with the consistency models was substantially faster, requiring only a handful of network evaluations compared to more than 100 for the diffusion-based approaches. Still, complete training and inference on 1,000 datasets required less wall-clock time than MCMC on a single dataset (under one day versus roughly three days).
The discrete consistency model matched the NRMSE of flow matching with a uniform noise schedule but showed a modest increase in calibration error, while the continuous consistency model exhibited weaker contraction and correspondingly higher calibration error. Flow matching with the power-law noise schedule performed slightly worse than the other flow matching variants but remained competitive overall. Using an exponential moving average (EMA) did not improve performance for this task (Section A.1).
@@
In high dimensional settings, the picture changes. Flow matching with the standard uniform noise schedule becomes competitive. Diffusion models continue to work well and again benefit from the VP noise schedule combined with $\mathbf{F}$- or $\mathbf{v}$-parameterization. In contrast, direct noise prediction is prone to instability and should be avoided. Discrete consistency models remain usable but tend to trade off accuracy for speed.
@@ 6.3 Case Study 3: Gaussian Random Fields as a Scalable SBI Benchmark
Gaussian random fields (GRFs) can model spatial uncertainty when the quantity of interest is a latent field rather than a small parameter vector (Figure 12C). Under the assumptions of homogeneity and isotropy, a GRF is fully described by its mean and covariance, or equivalently, by a power spectrum, allowing complex spatial patterns to be summarized by a few interpretable parameters, such as overall variance and correlation structure. This makes GRFs a standard prior in applications where spatially correlated properties are only indirectly observed, including subsurface permeability fields in hydrology [chen2025-grfgroundwater], heterogeneous material properties in computational geosciences [liu2019-grfreview], or the primordial matter distribution in cosmology [wandelt2013-grfcosmology]. In these domains, uncertainty quantification typically focuses on either low-dimensional parameters governing the spectrum of fluctuations or high-dimensional latent fields, often under low simulation budgets.
@@
We generate 2D GRFs with a power-law spectrum $P(k)\propto k^{-\alpha}\sigma$ using FyeldGenerator [cadiou2022-fyeldgenerator], and consider field resolutions $r\in\{8,16,32,64,128,256\}$. The parameters are drawn from Gaussian priors

### v2

@@ 6 Empirical Evaluation
A variety of design choices have been proposed for diffusion models. To assess their practical impact in SBI, we compare specific configurations and derive practical recommendations. For clarity, we adopt a fixed training protocol (Section A.1) and vary only the model family, the parameterization, or the noise schedule, while training models using the noise-prediction loss in (10). All models were trained using BayesFlow 2.0.8 [140].
@@ 6.1 Case Study 1: Low-Dimensional Benchmarks
Our first study presents the most comprehensive evaluation to date of diffusion models as posterior estimators on popular toy benchmark problems [114]. This study aims to highlight differences between the numerous design choices discussed in the preceding sections for the easiest types of problems (low-dimensional parameters, low-dimensional observations, easily obtainable high simulation budgets, see Table 2).
@@
  • Diffusion models under different noise schedules, using EDM noise schedules (variance-preserving (VP) or variance-exploding (VE)) with EDM weighting, and diffusion models using a cosine VP schedule with sigmoid or likelihood weighting;
@@
Figure 10: Comparison of inference performance across design choices on the benchmark by [114] (Case Study 1). We report mean and standard deviation of the C2ST, comparing samples from the diffusion model and the ground truth posterior samples for the 10 datasets from the benchmark (C2ST of 0.5 means approximate and reference posterior cannot be distinguished). As a baseline, we report the mean C2ST for the NPE approach in [114] trained on a budget of 100,000 simulations, showing how the difficulty across the problems differs.
Our analysis commences with the widely used set of ten benchmark models from [114]. These models span a range of low parameter and observation dimensionalities; for detailed descriptions, we refer to the original study by [114]. Because simulations are inexpensive, and all problems are low-dimensional, we train each model for 1,000 epochs, generating 30,000 new simulations per epoch. All neural networks are simple MLPs with residual layers and FiLM conditioning for the time without a separate summary network (Section A.1.1). To assess performance, we evaluate each trained model on the ten held-out simulated observations and compare the inferred posteriors with the benchmark “ground-truth” posteriors. We report the classifier two-sample test (C2ST) for a direct comparison with benchmark baselines (see Section A.1 for a definition).
@@
Diffusion models with a variance-preserving EDM noise schedule achieved the best overall performance, closely followed by cosine-based schedules (Figure 10). In contrast, the variance-exploding EDM schedule performed worse on several tasks, often producing higher C2ST scores. In terms of weighting, EDM seems to have a slight edge over sigmoid and likelihood weighting (Figure A.1).
@@
Across tasks, differences among the $\mathbf{F}$-, $\boldsymbol{v}$-, and $\boldsymbol{\epsilon}$-parameterizations for the respective optimal sampler were generally small (Figure 10). However, the $\boldsymbol{\epsilon}$-parameterization seems to be much more sensitive to the choice of the sampler (Figure A.1). EMA smoothing did not yield systematic improvements. In most cases, EMA and non-EMA variants performed similarly.
@@
Flow matching models were competitive on all benchmarks but showed a slight overall performance drop on the smallest problems compared to diffusion models (Figure 10). Flow matching and the conditional optimal-transport (COT) variant performed comparably with no clear trend, whereas ignoring the condition for computing the OT plan seems to harm accuracy (Figure A.2). Similar to previous studies [190, 128], the power-law noise schedule with $\rho=-0.6$ improved performance in most tasks, especially for the more difficult Lotka-Volterra task.
@@
Across diffusion and flow matching models, the adaptive two-step SDE solver was both the fastest and the most accurate (Figure 10). Performance differences among ODE and SDE solvers were small, except for the $\boldsymbol{\epsilon}$-parameterization, which shows more variability (Figure A.1). However, annealed Langevin sampling did not reach the accuracy of ODE or SDE solvers, even with 4 times as many steps, and the predictor–corrector scheme also did not improve C2ST scores (Figure A.1). For ODEs, adaptive higher-order methods consistently outperformed Euler: they were faster and achieved lower error (but with only minor differences) by taking larger, fewer steps, with TSIT5 showing a slight advantage over RK45. Step rejection rates in the adaptive methods aligned with trajectory smoothness: flow matching trajectories did not trigger rejections; EDM-based diffusion models triggered fewer than ten; and cosine-schedule models typically caused ten to twenty.
@@
For low-dimensional problems, we recommend to use diffusion models with a variance-preserving EDM noise schedule with $\mathbf{F}$ parameterization. Across problems and samplers, this emerges as the most reliable choice. For some problems that require higher contraction in the posterior, a variance-exploding noise schedule might be nevertheless beneficial. Moreover, we recommend the usage of adaptive solvers and, in particular, the two-step adaptive SDE solver in light of faster and more accurate inference. Only the consistency models are faster at the cost of potentially reduced accuracy. We discourage the use of annealed Langevin sampling or predictor-corrector sampling, as they are harder to tune and can lead to poor results.
@@ 6.2 Case Study 2: Large-Scale ODE Benchmark
Figure 11: Case Study 2: Comparison of inference performance across design choices. A Visualization of the reactions described by the [13] model. The parameters can vary per experimental conditions and some also for repeated experiments, leading to 72 parameters to be estimated. Blue indicates observed states. B Model simulation (with noise). C Best performing model from each family. We also show MCMC as a baseline. All metrics were scaled such that the optimal value would be 1 (only contraction has no optimal value) (Section A.1). D Diffusion models with cosine noise schedules and different parameterizations: $\mathbf{F}$, $\boldsymbol{v}$, and $\boldsymbol{\epsilon}$ using the ODE sampler. E Diffusion models with variance-preserving and exploding EDM noise schedules using the ODE sampler.
@@
To evaluate the capabilities of diffusion models for SBI in actual practice, we consider a mechanistic model from the PEtab collection of ODE models [67]. As a representative high-dimensional benchmark, we consider the [13] model, which represents a quantitative description of the enzymatic reactions involved in indigoidine biosynthesis (Figure 11A). The model includes experimental condition specific parameters for bacterial growth ($\beta$, $\text{Bac}_{\text{max}}$ and time lag $\tau$), conversion of glutamine (Gln) to cyclized glutamine (cGln) ($k_{\text{syn}}$), dimerization of cGln to indigoidine (Ind) ($k_{\text{dim}}$) and degradation of indigoidine ($k_{\text{deg}}$). The model comprises $19$ experimental conditions with more than $\text{dim}(\mathbf{y})=27{,}000$ measurements (Figure 11B) and $\text{dim}(\boldsymbol{\theta})=72$ parameters. It is one of the most data-rich ODE models in the collection, making it particularly suitable for testing the capacity of inference algorithms to handle large parameter spaces (see Section A.1.2 for more details).
@@
Model performance was assessed on 1,000 datasets using scaled calibration error (scaled such that its maximum is 1), contraction, normalized root mean squared error (NRMSE), and C2ST scores. For C2ST, the classifier is trained on posterior samples together with the log-likelihood as a data summary and compared to samples from the prior [193]. For a single scalar ranking across methods, we report the sum of empirical calibration error (ECE) and NRMSE (see Section A.1 for a definition of the metrics).
@@
Flow Matching OT ODE 0.116 (0.002) 0.033 (0.013) 0.734 (0.024) 0.602 1
Uniform ODE 0.116 (0.002) 0.037 (0.015) 0.733 (0.024) 0.615 2
$\rho=-0.6$ ODE 0.123 (0.003) 0.035 (0.014) 0.689 (0.023) 0.586 5
Diffusion Cosine, VP, $\mathbf{v}$ ODE 0.123 (0.003) 0.033 (0.013) 0.684 (0.024) 0.575 3
Cosine, VP, $\mathbf{F}$ ODE 0.122 (0.003) 0.034 (0.014) 0.696 (0.021) 0.596 4
Cosine, VP, $\mathbf{F}$ SDE 0.119 (0.002) 0.040 (0.017) 0.706 (0.021) 0.601 6
Cosine, VP, $\boldsymbol{\epsilon}$ SDE 0.119 (0.003) 0.040 (0.018) 0.704 (0.033) 0.603 7
Cosine, VP, $\mathbf{v}$ SDE 0.122 (0.002) 0.041 (0.016) 0.686 (0.022) 0.577 8
EDM, VP, $\mathbf{F}$ ODE 0.127 (0.003) 0.036 (0.014) 0.670 (0.030) 0.587 9
EDM, VP, $\mathbf{F}$ SDE 0.123 (0.002) 0.045 (0.018) 0.686 (0.029) 0.567 10
EDM, VE, $\mathbf{F}$ SDE 0.179 (0.003) 0.093 (0.034) 0.411 (0.055) 0.649 14
EDM, VE, $\mathbf{F}$ ODE 0.128 (0.002) 0.157 (0.055) 0.807 (0.027) 0.762 15
Cosine, VP, $\boldsymbol{\epsilon}$ ODE 0.281 (0.014) 0.069 (0.027) 0.005 (0.004) 0.536 17
Diffusion (EMA) EDM, VP, $\mathbf{F}$ ODE 0.128 (0.003) 0.044 (0.019) 0.667 (0.032) 0.551 11
EDM, VP, $\mathbf{F}$ SDE 0.124 (0.002) 0.056 (0.024) 0.683 (0.028) 0.570 12
Consistency Discrete ODE 0.120 (0.002) 0.073 (0.035) 0.705 (0.027) 0.562 13
Continuous ODE 0.158 (0.004) 0.161 (0.067) 0.479 (0.055) 0.529 16
MCMC - MCMC 0.124 (0.000) 0.427 (0.217) 1.000 (0.000) 0.889 18
@@ Results
Across both summary networks, several design choices consistently led to improved accuracy and robustness (Figure 11). Flow matching with a uniform noise schedule and its optimal transport variant achieved the strongest overall performance, exhibiting low error and stable behavior. Diffusion models with cosine schedules in the $\mathbf{v}$- and $\mathbf{F}$-parameterizations—using either ODE or SDE samplers—performed similarly well. The VP EDM schedule also yielded competitive accuracy. In contrast, the VE EDM schedule produced higher contraction and NRMSE and suffered from increased calibration error in this setting (Table 4), aligning with observations from the low-dimensional benchmarks. Moreover, C2ST showed lower sensitivity in this high-dimensional experiment, even when using log-likelihood summaries, suggesting that it does not reliably discriminate between posterior approximations in this regime (Table 4). In particular, when the update from the prior to the posterior was small, as for $\boldsymbol{\epsilon}$-prediction models, C2ST alone is unreliable as an accuracy measure. The relative ranking of inference methods was consistent across both summary networks, indicating that the performance differences are driven by the inference algorithms themselves rather than the choice of summary backbone.
Training times were similar across most methods, only flow matching OT took twice as long due to the added computation of optimal couplings; few-step inference with the consistency models was substantially faster, requiring only a handful of network evaluations compared to more than 100 for the diffusion-based approaches. Still, complete training and inference on 1,000 datasets required less wall-clock time than MCMC on a single dataset (under one day versus roughly three days).
The discrete consistency model nearly matched the NRMSE of flow matching (OT, Uniform) but showed a modest increase in calibration error, while the continuous consistency model exhibited weaker contraction and correspondingly higher calibration error. Flow matching with the power-law noise schedule performed slightly worse than the other flow matching variants but remained competitive overall. Using an exponential moving average (EMA) did not improve performance for this task (Section A.1).
@@
In high dimensional settings, the picture changes. Flow matching with the standard uniform noise schedule becomes competitive. Diffusion models continue to work well and again benefit from the VP noise schedule combined with $\mathbf{F}$- or $\mathbf{v}$-parameterization, while the cosine schedule has a very slight advantage compared to EDM. In contrast, direct noise prediction is prone to instability and should be avoided. Discrete consistency models remain usable but tend to trade off accuracy for speed.
@@ 6.3 Case Study 3: Gaussian Random Fields as a Scalable SBI Benchmark
Gaussian random fields (GRFs) can model spatial uncertainty when the quantity of interest is a latent field rather than a small parameter vector (Figure 12C). Under the assumptions of homogeneity and isotropy, a GRF is fully described by its mean and covariance, or equivalently, by a power spectrum, allowing complex spatial patterns to be summarized by a few interpretable parameters, such as overall variance and correlation structure. This makes GRFs a standard prior in applications where spatially correlated properties are only indirectly observed, including subsurface permeability fields in hydrology [21], heterogeneous material properties in computational geosciences [108], or the primordial matter distribution in cosmology [186]. In these domains, uncertainty quantification typically focuses on either low-dimensional parameters governing the spectrum of fluctuations or high-dimensional latent fields, often under low simulation budgets.
@@
We generate 2D GRFs with a power-law spectrum $P(k)\propto k^{-\alpha}\sigma$ using FyeldGenerator [19], and consider field resolutions $r\in\{8,16,32,64,128,256\}$. The parameters are drawn from Gaussian priors